Find the Laplace transforms of the following function:

8. L{t^2 e^-2t + e^-t cos2t + 3}


1
Expert's answer
2022-05-24T17:58:26-0400
"F(s)=L(t^2e^{-2t}+e^{-t}\cos(2t)+3)"

"=\displaystyle\int_{0}^{\infin}e^{-st}(t^2e^{-2t}+e^{-t}\cos(2t)+3)dt"

"\int t^2e^{-(s+2)t}dt=-\dfrac{t^2e^{-(s+2)t}}{s+2}+\dfrac{2}{s+2}\int te^{-(s+2)t}dt"

"=-\dfrac{t^2e^{-(s+2)t}}{s+2}-\dfrac{2te^{-(s+2)t}}{(s+2)^2}"


"+\dfrac{2}{(s+2)^2}\int e^{-(s+2)t}dt"

"=-\dfrac{t^2e^{-(s+2)t}}{s+2}-\dfrac{2te^{-(s+2)t}}{(s+2)^2}-\dfrac{2e^{-(s+2)t}}{(s+2)^3}+C"


"L(t^2e^{-2t})=\lim\limits_{A\to\infin}\displaystyle\int_{0}^{\infin}e^{-st}(t^2e^{-2t})dt"

"=[-\dfrac{t^2e^{-(s+2)t}}{s+2}-\dfrac{2te^{-(s+2)t}}{(s+2)^2}-\dfrac{2e^{-(s+2)t}}{(s+2)^3}]\begin{matrix}
 A \\
 0
\end{matrix}"

"=-0-0-0-(-0-0-\dfrac{2}{(s+2)^3})=\dfrac{2}{(s+2)^3}"

"\int e^{-(s+1)t}\cos(2t)dt=-\dfrac{e^{-(s+1)t}\cos(2t)}{s+1}"

"-\dfrac{2}{s+1}\int e^{-(s+1)t}\sin (2t)dt=-\dfrac{e^{-(s+1)t}\cos(2t)}{s+1}"

"+\dfrac{2e^{-(s+1)t}\sin(2t)}{(s+1)^2}-\dfrac{4}{(s+1)^2}\int e^{-(s+1)t}\cos (2t)dt"

"((s+1)^2+4)\int e^{-(s+1)t}\cos (2t)dt"

"=-e^{-(s+1)t}\cos(2t)(s+1)+2e^{-(s+1)t}\sin(2t)"

"\int e^{-(s+1)t}\cos (2t)dt=-\dfrac{e^{-(s+1)t}\cos(2t)(s+1)}{s^2+2s+5}"

"+\dfrac{2e^{-(s+1)t}\sin(2t)}{s^2+s+5}+C_1"


"L(e^{-t}\cos(2t))=\lim\limits_{A\to\infin}\displaystyle\int_{0}^{\infin}e^{-st}(e^{-t}\cos(2t))dt"

"=[-\dfrac{e^{-(s+1)t}\cos(2t)(s+1)}{s^2+2s+5}+\dfrac{2e^{-(s+1)t}\sin(2t)}{s^2+s+5}]\begin{matrix}
 A \\
 0
\end{matrix}"

"=-0+0-(-\dfrac{s+1}{s^2+2s+5}+0)=\dfrac{s+1}{s^2+2s+5}"



"L(3)=\lim\limits_{A\to\infin}\displaystyle\int_{0}^{\infin}e^{-st}(3)dt"

"=[-\dfrac{3e^{-st}}{s}+]\begin{matrix}
 A \\
 0
\end{matrix}=-0+\dfrac{3}{s}=\dfrac{3}{s}"


"L(t^2e^{-2t}+e^{-t}\cos(2t)+3)"

"=\dfrac{2}{(s+2)^3}+\dfrac{s+1}{s^2+2s+5}+\dfrac{3}{s}"